# Coarse obstructions to cocompact cubulation

### v2

@@ Abstract.
We provide geometric methods to both lower- and upper-bound the large-scale dimension of CAT(0) cube complexes quasiisometric to a given group $G$. When these bounds overlap, this provides obstructions to $G$ being cocompactly cubulated. More strongly, it prevents $G$ from being a coarse median space.
As applications, we show that many free-by-cyclic groups cannot be cocompactly cubulated, and that any tubular group with a coarse median is virtually compact special.
@@ 1. Introduction
Many groups have been successfully (cocompactly) cubulated, and many of these are even virtually special; indeed, Agol’s theorem [Ago13] states that every hyperbolic group that can be cocompactly cubulated is virtually special. However, there are many groups of interest for which cocompact cubulability is unknown. For many, it is expected to be impossible; we simply lack refined criteria to verify this suspicion. In fact, all general methods for obstructing cocompact cubulation known to the authors are simply negations of properties of cocompactly cubulated groups. For instance, if $G$ has super-quadratic Dehn function, property (T), distorted elements, or elements whose centralisers do not virtually split, then $G$ is not cocompactly cubulated [NR97, Ebe82].
@@
If $G$ acts properly cocompactly on a CAT(0) cube complex $X$, then $G$ is quasiisometric to $X$. Merely being quasiisometric to a CAT(0) cube complex (quasicubical) is much weaker than being cocompactly cubulated; for instance, all hyperbolic groups are quasicubical [HW12], even though some have property (T) and hence cannot act on CAT(0) cube complexes without global fixed-points [NR97]. If $G$ is quasicubical then one can pass the median of the cube complex along the quasiisometry to equip $G$ with a ternary operator that behaves coarsely the same. This makes $G$ a coarse median space in the sense of [Bow13].
@@
Let $G$ be a finitely generated group with $\operatorname{qf.rk}G\leq n$. If $G$ contains a quasiisometrically embedded richly branching flat of dimension $n$, then $G$ cannot be a coarse median space. In particular, $G$ cannot be cocompactly cubulated.
@@
As the name suggests, then, richly branching flats should be thought of as being flats with “too much” branching “all over”. Similar configurations were considered in work of Haettel [Hae16], who characterised which symmetric spaces and affine buildings admit coarse medians. However, his arguments rely on work specific to that setting, such as [KL97].
@@
Part of the proof of Theorem A involves relating the quasiflat rank of a coarse median space $X$ to its coarse-median rank, $\operatorname{rk}X$. Proposition 3.1 in particular shows that these two quantities agree when $X$ is a finitely generated group. In general, though, it is not easy to ascertain the quasiflat rank of a group. In order to facilitate the applications of Theorem A discussed below, we therefore establish the following bound, which is a consequence of Proposition 3.1 and the general Lemma 3.4. Recall that the virtual geometric dimension $\operatorname{vgd}G$ of a group $G$ is the infimal dimension among classifying spaces of finite-index subgroups of $G$.
@@
Although this bound is far from optimal in general, as can be seen from hyperbolic manifolds, it can be useful in concrete situations where geometric (or cohomological) dimensions are easily computable.
@@
It is therefore natural to ask which free-by-cyclic groups can be cocompactly cubulated. As it turns out, all hyperbolic free-by-cyclic groups are cocompactly cubulated, by work of Hagen–Wise [HW16, HW15]. (More generally, all hyperbolic hyperbolic-by-cyclic groups are [DKM23].) Also, Hagen–Przytycki characterised which graph manifold groups are cocompactly cubulated [HP15], and some of those are free-by-cyclic. Gersten’s group [Ger94a] (and tweaks thereof) does not act properly on a $\mathrm{CAT}(0)$ space, for reasons involving translation-length, and so cannot be cocompactly cubulated. Though cocompact cubulability is expected to be rare amongst non-hyperbolic free-by-cyclic groups, little seems to be known in general. For example, the following question is open.
@@
If a free-by-cyclic group $G$ (virtually) is represented by an improved relative train track such that some Nielsen cycle either
  • supports at least three internal linear strata, or
  • supports two internal linear strata and has a nearby source,
then $G$ cannot admit a coarse median.
The way to think about Theorem C is as follows. The improved relative train track structure (from [BFH00]; see Item 5.1) provides a systematic way to build up $G=F\rtimes_{\phi}\mathbf{Z}$ in layers known as strata. The simplest strata are those that are fixed by $\phi$; these can give rise to $\mathbf{Z}^{2}$ subgroups of $G$. More generally, Nielsen cycles are loops in the train track that are fixed. Next simplest are the linear strata: roughly, the translates of a linear stratum by powers of $\phi$ grow in complexity at a linear rate. These can give rise to quasiflats that branch off the above $\mathbf{Z}^{2}$ subgroups. The terms “internal” and “source” ensure that this happens (Definition 5.5). The conditions of Theorem C can therefore be understood as saying that there is some quasiflat in $G$ with three directions of branching, which allows the obstructions of Theorem A to kick in.
Let us say that a free-by-cyclic group has rich linearity if it satisfies the assumptions of Theorem C. The class of free-by-cyclic groups with rich linearity is larger than those for which Gersten’s ideas rule out being CAT(0); a simple example is provided in Example 5. Gersten’s group itself fits into the second case of the theorem.
@@
If a free-by-cyclic group $G$ is represented by an improved relative train track with no quadratic strata and such that each Nielsen cycle supports at most one linear stratum, then $G$ is quasicubical.
In the special case of linearly-growing free-by-cyclic groups (or, more generally, those with no quadratic strata), the combination of Theorems C and D leaves quite a restricted set of possible train tracks. It would be interesting to know whether such groups are quasicubical. The following is related to a possible converse to Theorem C; see Remark 5 for more discussion and related questions. Note that such a converse would imply that rich linearity is witnessed by all fibrations.
@@
If a free-by-cyclic group has no quasiisometrically embedded richly branching flats, must it be quasiisometric to a finite-dimensional CAT(0) cube complex?
@@
In [Wis14b], Wise characterised which tubular groups admit free actions on CAT(0) cube complexes. These cubulations were further investigated by Woodhouse [Woo16, Woo18]. Wise also showed that all cocompactly cubulated tubular groups are virtually special.
@@
We observe that this result of Haettel answers negatively a question asked by Wise in the 2014 ICM proceedings [Wis14a, Prob. 13.40].
@@
We are indebted to Naomi Andrew for several very useful discussions about free-by-cyclic groups. We are very grateful to Mark Hagen and Monika Kudlinska for helpful conversations about train tracks. We thank Jason Behrstock and Jean Pierre Mutanguha for comments on an earlier version. We also thank the organisers of the 2023 thematic program on geometric group theory in Montreal, where this work began.
@@ 3. Dimension bounds
Whilst Proposition 3.1 is very precise, it is not always easy to ascertain the quasiflat rank of a group. It is therefore desirable to have a result giving less optimal but more pragmatic control. That is the goal of the remainder of this section, where we give a bound in terms of geometric dimension.
@@
For a finitely generated group $G$, write $\operatorname{gd}G$ for the geometric dimension of $G$; that is, the infimal dimension of a $K(G,1)$. Recall that $G$ has type $F$ it has a finite $K(G,1)$, and type $F_{\infty}$ if it has a $K(G,1)$ whose $n$–skeleton is finite for all $n$.
3.4 Lemma.
@@
For a finitely generated group $G$, write $\operatorname{vgd}G$ for the virtual geometric dimension of $G$; that is, the infimal geometric dimension of a finite-index subgroup of $G$.
@@
By Proposition 3.1, we have $\operatorname{rk}G=\operatorname{qf.rk}G$. According to [Bow22, Prop. 12.4.7], every asymptotic cone of $G$ is $n$–connected for all $n$. By [Ril03, Thm D], we find that $G$ is of type $F_{\infty}$. Suppose that $\operatorname{vgd}G<\infty$, and let $H$ be a finite-index subgroup of $G$ realising $\operatorname{vgd}G$. It is of type $F_{\infty}$, for instance by [Bro94, IX.6.1]. The result follows from Lemma 3.4, because $H$ is quasiisometric to $G$. ∎
3.6 Remark.
Since coarse medians can be passed along quasiisometries, Theorem 3.5 could equally well be stated in terms of the infimal geometric dimension of a group quasiisometric to $G$.
@@ 5. Free-by-cyclic groups
Note that no free factors of $F$ are fixed by any power of $\phi$, so for any IRTT structure on a power of $\phi$ the invariant strata can only form a subforest.
@@
We now turn to the case where $\phi$ has more linear strata. Our goal will be to show that the equivalent of Proposition 5.2 fails in a strong way for many such $\phi$, by finding RBFs. We shall need the following result, which seems to be known to experts. We are grateful to Monika Kudlinska for informing us of it, and to Jean Pierre Mutanguha for sharing his preprint [Mut] with us. Recall that all free groups considered here are finitely generated.
5.3 Proposition ([Mut, Lem. 4.1, 4.2]).
@@
As $G$ has geometric dimension two, Theorem 3.5 shows that if $G$ admits a coarse median, then it admits one of rank at most two. Our goal is to find a 2–RBF in $G$, for then Theorem 4.3 will show that $G$ can admit no such coarse median. There is no loss in replacing $\phi$ by a positive power, because this replaces $G$ by a finite-index subgroup. We can therefore let $G=F\rtimes_{\phi}\langle t\rangle$ witness rich linearity.
@@
All quasicubical metric spaces have coarse medians, but the converse is not true in general. However, in the setting of finitely presented groups, examples that distinguish the classes are sorely lacking. In many cases, a coarse median can actually be promoted to quasicubicality [HP22, Pet21]. We expect that free-by-cyclic groups do not distinguish the classes, which explains the phrasing of Question 2.
@@ 6.2. Coarse median tubular groups
Let $G$ be a rank-$2$ tubular group. If $G$ admits a coarse median, then $G$ is cocompactly cubulated and virtually compact special.
@@
According to [Bow13, Cor. 8.3], if $G$ admits a coarse median then it has a quadratic isoperimetric function, and so Theorem 6.8 shows that $G$ must be undistorted. Moreover, $G$ has geometric dimension two since it is a graph of two dimensional groups (see [SW79, Prop. 3.6]), so Theorem 3.5 shows that $\operatorname{rk}G\leq 2$. In particular, Theorem 4.3 implies that $G$ cannot have a quasiisometrically embedded 2–RBF. By Lemma 6.9, this means that no vertex group of $G$ can have more than two commensurability classes of incident edge groups. Since $G$ is undistorted, it contains no Baumslag–Solitar subgroups $\operatorname{BS}(m,n)$ with $m\neq\pm n$. Consequently, [Wis14b, Cor 5.10, 5.9] implies that $G$ is cocompactly cubulated and virtually compact special. ∎
By a result of Woodhouse [Woo18], it follows that, for tubular groups, admitting a coarse median is equivalent to the existence of a free action on a locally finite CAT(0) cube complex. Using Wise’s characterisation [Wis14b], there are tubular groups that are freely cubulable but do not admit coarse medians.
6.11 Example (A CAT(0), freely cubular, tubular, non-quasicubular group).
Let $T$ be the $1$–skeleton of a tetrahedron, and 3–edge-colour $T$ with colours $c_{h},c_{v},c_{d}$. Consider the tubular group $G$ on $T$ where $G_{e^{\pm}}$ is generated by: $(1,0)$ if $e$ has colour $c_{h}$; $(0,1)$ if $e$ has colour $c_{v}$; and $(1,1)$ if $e$ has colour $c_{d}$.
It is easy to see from Wise’s characterisation that $G$ acts freely on a $\mathrm{CAT}(0)$ cube complex. Moreover, the choice of edge-inclusions makes $X$ a CAT(0) space. However, since $G$ is undistorted and contains a vertex with three incident, non-commensurable edge groups, Lemma 6.9 implies $G$ has no coarse median (and in particular is not quasicubical). Groups such as $G$ can be viewed as “opposite” to hyperbolic groups with property (T), which are quasicubical yet do not act freely on any $\mathrm{CAT}(0)$ cube complex.
There is some overlap between Theorem 6.10 and Theorem 5.7. By combining [BD14, Cor. 4.17] and [Mac02, Thm 1.2], it can be seen that any tubular group that is free-by-cyclic is linearly growing. For such groups, one should then compare Proposition 5.2 and Theorem 5.7 with [Wis14b, Cor. 5.10].

### v3

@@ Abstract.
We provide geometric methods to give bounds on the large-scale dimension of CAT(0) cube complexes quasiisometric to a given group $G$. In situations where these bounds conflict we obtain obstructions to $G$ being cocompactly cubulated. More strongly, the obstructions prevent $G$ from being a coarse median space.
As applications, we show that many free-by-cyclic groups cannot be cocompactly cubulated, even virtually, and prove that any tubular group with a coarse median is virtually compact special. We also exhibit a group that is CAT(0), $C(6)$, and virtually special, yet is not quasiisometric to any CAT(0) cube complex. This is the first example of a $C(6)$ group that cannot be cocompactly cubulated, resolving a question of Jankiewicz and partially answering a question of Wise.
@@ 1. Introduction
Many groups have been successfully (cocompactly) cubulated, and many of these are even virtually special; indeed, Agol’s theorem [Ago13] states that every hyperbolic group that can be cocompactly cubulated is virtually compact special. However, there are many groups of interest for which cocompact cubulability is unknown. For many, it is expected to be impossible; we simply lack refined criteria to verify this suspicion. In fact, all general methods for obstructing cocompact cubulation known to the authors are simply negations of properties of cocompactly cubulated groups. For instance, if $G$ has super-quadratic Dehn function, property (T), distorted elements, or elements whose centralisers do not virtually split, then $G$ is not cocompactly cubulated [NR97, Ebe82].
@@
If $G$ acts properly cocompactly on a CAT(0) cube complex $X$, then $G$ is quasiisometric to $X$. Merely being quasiisometric to a CAT(0) cube complex (quasicubical) is much weaker than being cocompactly cubulated; for instance, all hyperbolic groups are quasicubical [HW12], even though some have property (T) and hence cannot act on CAT(0) cube complexes without global fixed-points [NR97]. If $G$ is quasicubical then one can pass the median of the cube complex along the quasiisometry to equip $G$ with a ternary operator that behaves coarsely the same. This makes $G$ a coarse median space in the sense of [Bow13]. Thus proving that a space does not admit any coarse median is much stronger than proving it cannot be cocompactly cubulated.
@@
Let $G$ be a finitely generated group with $\operatorname{qf.rk}G\leq n$. If $G$ contains a quasiisometrically embedded richly branching flat of dimension $n$, then $G$ cannot be a coarse median space. In particular, $G$ cannot be virtually cocompactly cubulated.
@@
As the name suggests, then, richly branching flats should be thought of as being flats with “too much” branching “all over”. Similar configurations were considered in work of Haettel [Hae16], who characterised which symmetric spaces and affine buildings admit coarse medians (see Theorem H). However, his arguments rely on work specific to that setting, such as [KL97].
@@
Part of the proof of Theorem A involves relating the quasiflat rank of a coarse median space $X$ to its coarse-median rank, $\operatorname{rk}X$. Proposition 3.1 in particular shows that these two quantities agree when $X$ is a finitely generated group. In general it is not easy to ascertain the quasiflat rank of a group, but we nevertheless have the following as a consequence of Proposition 3.1, which will be used in our applications of Theorem A below.
@@
Let $G$ be a finitely generated group. If $G$ admits a coarse median, then $\operatorname{rk}G\leq\operatorname{vcd}G$.
Although this bound is not optimal in general, as can be seen from hyperbolic manifolds, it can be useful in concrete situations where cohomological dimensions are easily computable.
@@
It is therefore natural to ask which free-by-cyclic groups can be cocompactly cubulated. As it turns out, all hyperbolic free-by-cyclic groups are cocompactly cubulated, by work of Hagen–Wise [HW16, HW15]. (More generally, all hyperbolic hyperbolic-by-cyclic groups are [DKM23].) Also, Hagen–Przytycki characterised which graph manifold groups are cocompactly cubulated [HP15], and some of those are free-by-cyclic. Gersten’s group [Ger94a] (and tweaks thereof) does not act properly on a $\mathrm{CAT}(0)$ space, for reasons involving translation-length, and so cannot be cocompactly cubulated. There are also examples that can be cubulated, but not cocompactly [WY25]. Though cocompact cubulability is expected to be rare amongst non-hyperbolic free-by-cyclic groups, little seems to be known in general. For example, the following question is open.
@@
If a free-by-cyclic group has rich linearity, then it cannot admit a coarse median.
The rough idea of the rich linearity condition (see Definition 5.6) in Theorem C is as follows. The free-by-cyclic group $G=F\rtimes_{\phi}\mathbf{Z}$ virtually has an improved relative train track structure (from [BFH00]; see Item 5.1), which provides a systematic way to build up $G$ in layers known as strata. The simplest strata are those that are fixed by $\phi$; these can give rise to $\mathbf{Z}^{2}$ subgroups of $G$. The translates of other strata by powers of $\phi$ grow in complexity. When this growth is linear, this can give rise to quasiflats that branch off the above $\mathbf{Z}^{2}$ subgroups. The rich linearity condition enforces that this happens enough to find a richly branching flat and employ Theorem A.
The class of free-by-cyclic groups with rich linearity is larger than those for which Gersten’s ideas rule out being CAT(0); a simple example that can be shown to be CAT(0) is provided in Example 5.
@@
If a free-by-cyclic group $G$ is represented by an improved relative train track with no quadratically growing strata and such that each Nielsen cycle supports at most one linear stratum, then $G$ is virtually a colourable hierarchically hyperbolic group, and in particular is quasicubical.
For free-by-cyclic groups of linear growth, the combination of Theorems C and D leaves quite a restricted set of possible train tracks. The following is related to a possible converse to Theorem C; see Remark 5 for more discussion and related questions. Note that such a converse would imply that rich linearity is witnessed by all fibrations.
@@
Is it true that a free-by-cyclic group is virtually hierarchically hyperbolic if and only if it does not have rich linearity? Less strictly, if a free-by-cyclic group has no quasiisometrically embedded richly branching flats, must it be quasiisometric to a finite-dimensional CAT(0) cube complex?
@@
In [Wis14b], Wise characterised which tubular groups admit free actions on CAT(0) cube complexes. These cubulations were further investigated by Woodhouse [Woo16, Woo18]. Wise also showed that all cocompactly cubulated tubular groups are virtually compact special.
As an instance of this characterisation, we construct a tubular group that has a $C(6)$ structure but does not admit any coarse median (Example 6.11). This is the first example of a $C(6)$ group that does not act properly cocompactly on any $\mathrm{CAT}(0)$ cube complex, answering a question asked by Jankiewicz [Jan18, Q. 6.6.4] and suggested by Wise [Wis04].
Theorem F.
There exists a group $G$ that is $C(6)$, $\mathrm{CAT}(0)$, and virtually special, yet does not admit any coarse median. In particular, $G$ is not virtually cocompactly cubulated.
Theorem F contrasts with the situation for $C^{\prime}(\frac{1}{6})$ groups, which are cocompactly cubulated [Wis04]. It is still open whether there is some $n$ for which all $C(n)$ groups are cocompactly cubulated. It also contrasts with the strict $C(6)$ condition introduced in [MW25]: the authors show that strict $C(6)$ groups are hyperbolic relative to virtually-$\mathbf{Z}^{2}$ subgroups, and they are therefore quasicubical.
@@
We observe that this result of Haettel answers negatively a question asked by Wise in the 2014 ICM proceedings [Wis14a, Prob. 13.40], because it provides CAT(0) spaces with proper cocompact group actions that are not quasiisometric to CAT(0) cube complexes. Theorem F provides a new example, and indeed one can use Theorem E to produce many such examples.
@@
We are indebted to Naomi Andrew for several very useful discussions about free-by-cyclic groups. We are very grateful to Mark Hagen and Monika Kudlinska for helpful conversations about train tracks. We thank Jason Behrstock, Pritam Ghosh, Rob Kropholler, Jean Pierre Mutanguha, and Xiaolei Wu for comments on an earlier version. We also thank the organisers of the 2023 thematic program on geometric group theory in Montreal, where this work began.
@@ 3. Dimension bounds
3.2 Remark (Asymptotic rank).
The behaviour in Example 3 arises because asymptotic cones are taken with a fixed basepoint and the space is not homogeneous. However, there is still something to be said even without the assumptions of properness and cocompactness. Following [Wen11] (see Proposition 3.1 thereof), the asymptotic rank of a metric space $X$ can be defined to be the supremal $n$ such that there is some asymptotic cone (with basepoints allowed to move) $\hat{X}$ of $X$ and a sequence of subspaces $B_{k}\subset X$ whose limit in $\hat{X}$ is the unit ball in some normed space $(\mathbf{R}^{n},\|\cdot\|)$.
The same proof as in Proposition 3.1 shows that if $X$ is a coarsely connected coarse median space of rank $d$, then the asymptotic rank of $X$ is $d$. Indeed, consider the subset $Q_{n}$ constructed in the proof of Proposition 3.1, which is the image of a product of $d$ real intervals of length $n$ under a uniform-quality quasimedian quasiisometric embedding. Let $z_{n}$ be the centre of $Q_{n}$. The limit of the sequence $(Q_{n})$ in the asymptotic cone $\hat{X}=\lim_{\omega}(X,\frac{1}{\sqrt{n}},z_{n})$ is $(\mathbf{R}^{d},\ell^{1})$, which implies the existence of the desired sequence. The cocompactness was only used to control the basepoint, and the properness was only used to say that the space was equal to its own ultralimit.
Whilst Proposition 3.1 is very precise, it is not always easy to ascertain the quasiflat rank of a given group $G$. In practice, therefore, it is useful to have a statement in terms of a more easily calculable quantity. The following is Theorem B from the introduction.
3.3 Corollary.
Let $G$ be a finitely generated group. If $G$ admits a coarse median, then $\operatorname{rk}G\leq\operatorname{vcd}G$.
Proof.
According to [Sau06, Thm 1.2], any group that contains a quasiflat of dimension $n$ must have cohomological dimension at least $n$. Since Proposition 3.1 shows that $\operatorname{rk}G=\operatorname{qf.rk}G$, this proves the result. ∎
By the Eilenberg–Ganea theorem, if a finitely presented group $G$ has cohomological dimension not equal to two, then its cohomological dimension is equal to its geometric dimension: the infimal dimension of a $K(G,1)$. Moreover, it remains unknown whether the same holds for groups of cohomological dimension two. Thus, in almost all cases Corollary 3.3 can be equivalently stated using virtual geometric dimension. In this case the application of Sauer’s theorem, whose statement is considerably more general than is being used here, can be replaced by Proposition 3.6 below, which has a comparatively short proof.
@@
For a finitely generated group $G$, write $\operatorname{gd}G$ for the geometric dimension of $G$. Recall that $G$ has type $F$ it has a finite $K(G,1)$, and type $F_{\infty}$ if it has a $K(G,1)$ whose $n$–skeleton is finite for all $n$.
3.6 Proposition.
@@
We conclude by proving a geometric-dimension variation of Corollary 3.3.
@@
By Proposition 3.1, we have $\operatorname{rk}G=\operatorname{qf.rk}G$. According to [Bow22, Prop. 12.4.7], every asymptotic cone of $G$ is $n$–connected for all $n$. By [Ril03, Thm D], we find that $G$ is of type $F_{\infty}$. Suppose that $\operatorname{vgd}G<\infty$, and let $H$ be a finite-index subgroup of $G$ realising $\operatorname{vgd}G$. It is of type $F_{\infty}$, for instance by [Bro94, IX.6.1]. The result follows from Proposition 3.6, because $H$ is quasiisometric to $G$. ∎
@@ 5. Free-by-cyclic groups
Note that no cyclic free factors of $F$ are fixed by any power of $\phi$, so for any IRTT structure on a power of $\phi$ the invariant strata can only form a subforest.
@@
We now turn to the case where $\phi$ has more linear strata. Our goal will be to show that the equivalent of Proposition 5.2 fails in a strong way for many such $\phi$, by finding RBFs. We shall need the following result, which seems to be known to experts. We are grateful to Monika Kudlinska for informing us of it, and to Jean Pierre Mutanguha for sharing his preprint [Mut24] with us. Recall that all free groups considered here are finitely generated.
5.3 Proposition ([Mut24, Lem. 4.1, 4.2]).
@@
As $G$ has geometric dimension two, Corollary 3.3 shows that if $G$ admits a coarse median, then it admits one of rank at most two. Our goal is to find a 2–RBF in $G$, for then Theorem 4.3 will show that $G$ can admit no such coarse median. There is no loss in replacing $\phi$ by a positive power, because this replaces $G$ by a finite-index subgroup. We can therefore let $G=F\rtimes_{\phi}\langle t\rangle$ witness rich linearity.
@@
All quasicubical metric spaces have coarse medians, but the converse is not true in general. However, in the setting of finitely generated groups, examples that distinguish the classes are sorely lacking. In many cases, a coarse median can actually be promoted to quasicubicality [HP22, Pet21]. We expect that free-by-cyclic groups do not distinguish the classes, which explains the phrasing of Question 2.
@@ 6.2. Coarse median tubular groups
Let $G$ be a tubular group. If $G$ admits a coarse median, then $G$ is cocompactly cubulated and virtually compact special.
@@
According to [Bow13, Cor. 8.3], if $G$ admits a coarse median then it has a quadratic isoperimetric function, and so Theorem 6.8 shows that $G$ must be undistorted. Moreover, $G$ has geometric dimension two since it is a graph of two dimensional groups (see [SW79, Prop. 3.6]), so Corollary 3.3 shows that $\operatorname{rk}G\leq 2$. In particular, Theorem 4.3 implies that $G$ cannot have a quasiisometrically embedded 2–RBF. By Lemma 6.9, this means that no vertex group of $G$ can have more than two commensurability classes of incident edge groups. Since $G$ is undistorted, it contains no Baumslag–Solitar subgroups $\operatorname{BS}(m,n)$ with $m\neq\pm n$. Consequently, [Wis14b, Cor 5.10, 5.9] implies that $G$ is cocompactly cubulated and virtually compact special. ∎
Using Wise’s characterisation [Wis14b], there are tubular groups that are freely cubulable but do not admit coarse medians.
6.11 Example (A freely cubular, tubular, non-quasicubular group).
Let $T$ be the $1$–skeleton of a tetrahedron, and 3–edge-colour $T$ with colours $g,p,b$. Up to an isomorphism of $T$, there is one way to do this. Consider the tubular group $G$ on $T$ where $G_{e^{\pm}}$ is generated by: $(1,0)$ if $e$ has colour $g$; $(0,1)$ if $e$ has colour $p$; and $(1,-1)$ if $e$ has colour $b$.
It is easy to see from Wise’s characterisation that $G$ acts freely on a $\mathrm{CAT}(0)$ cube complex. From work of Woodhouse [Woo18], it follows that $G$ is virtually special.
Moreover, the choice of edge-inclusions makes $X$ a CAT(0) space. However, since $G$ is undistorted and contains a vertex with three incident, non-commensurable edge groups, Lemma 6.9 implies $G$ has no coarse median (and in particular is not quasicubical). Groups such as $G$ can be viewed as “opposite” to hyperbolic groups with property (T), which are quasicubical yet do not act freely on any $\mathrm{CAT}(0)$ cube complex.
It is additionally possible to give a $C(6)$ structure to the graph of spaces associated to $G$. Each vertex space is a torus formed from a four-by-four grid of hexagons by identifying appropriate boundary edges. We then subdivide the edges, turning each hexagon to a dodecagon. Each edge annulus is formed from four hexagons, with edges subdivided. The attaching maps at each vertex are indicated in Figure 8. It can be verified by inspection that the cell structure is $C(6)$.
Figure 8. A $C(6)$ structure on a torus and three incident edge groups. The torus is a quotient of the hexagonal plane by translations along the black arrows, with each edge subdivided. The four-by-four hexagonal grid is a fundamental domain. The green, purple, and blue edge groups can be identified with the subgroups $\langle(1,0)\rangle$, $\langle(0,1)\rangle$, and $\langle(1,-1)\rangle$, respectively. The attaching maps send the bold vertices to bold vertices.
There is some overlap between Theorem 6.10 and Theorem 5.7. By combining [BD14, Cor. 4.17] and [Mac02, Thm 1.2], it can be seen that any tubular group that is free-by-cyclic is linearly growing. For such groups, one should then compare Proposition 5.2 and Theorem 5.7 with [Wis14b, Cor. 5.10]. Wise’s results and the interplay between free-by-cyclic and tubular groups have also been exploited in [WY25] to show that certain CAT(0) free-by-cyclic groups produced by Lyman [Lym23] are virtually special but not virtually cocompactly cubulated.